# Spec-Gloss Surfels and Normal–Diffuse Priors for Relightable Glossy Objects

### v1

@@ Abstract
Accurate reconstruction and relighting of glossy objects remain a longstanding challenge, as object shape, material properties, and illumination are inherently difficult to disentangle. Existing neural rendering approaches often rely on simplified BRDF models or parameterizations that couple diffuse and specular components, which restricts faithful material recovery and limits relighting fidelity. We propose a relightable framework that integrates a microfacet BRDF with the specular–glossiness parameterization into 2D Gaussian Splatting with deferred shading. This formulation enables more physically consistent material decomposition, while diffusion-based priors for surface normals and diffuse color guide early-stage optimization and mitigate ambiguity. A coarse-to-fine optimization of the environment map accelerates convergence and preserves high-dynamic-range specular reflections. Extensive experiments on complex, glossy scenes demonstrate that our method achieves high-quality geometry and material reconstruction, delivering substantially more realistic and consistent relighting under novel illumination compared to existing Gaussian splatting methods. The code is available at https://github.com/gkouros/SpecGloss-GS.
@@ 1 Introduction
Incorporating real-world objects into virtual environments is a fundamental task in applications such as digital twins and virtual reality. A major challenge arises in accurately reconstructing objects from the physical world and reproducing realistic lighting effects within specific virtual settings. This difficulty is particularly pronounced for highly reflective objects, where the lack of multi-view consistency renders traditional approaches, such as multi-view stereo (MVS), ineffective and makes reliable reconstruction itself a non-trivial problem.
@@
The choice of scene representation and rendering model introduces unique trade-offs in performance. For instance, approaches based on Radiance Fields [22, 26, 1] or 3D Gaussian Splatting (3DGS) [11] often fail to accurately capture object surfaces, thereby limiting their ability to model surface reflections. In contrast, methods employing Signed Distance Functions (SDFs) [32, 7, 21, 19] provide precise surface geometry, but their high computational cost poses challenges for real-time rendering. Another critical component of inverse rendering is the rendering model, which decomposes object materials into multiple attributes and employs predefined functions to synthesize appearance from target viewpoints. To achieve differentiability, various methods approximate the rendering equation integrals differently. For example, Ref-NeRF [30] uses an Integrated Directional Encoding, GaussianShader [9] employs specular GGX [31], Ref-GS [37] adopts a spherical-mip encoding, and 3DGS-DR [12] leverages deferred rendering. While these diverse formulations yield varying reconstruction quality, they share a fundamental limitation: the inverse solving process suffers from an inherent ambiguity, where multiple material-lighting combinations can produce identical appearances. Consequently, reconstructed material attributes and recovered lighting often deviate from ground truth values, leading to degraded rendering quality under novel lighting conditions. This is not a new observation, but a long-standing open problem going back to early work on intrinsic images [2], highlighted more recently in [14].
@@
In this work, we reduce ambiguity and improve relighting quality. We start by investigating the relationship between rendering models and relighting effects. Specifically, we analyze the coupling between object materials and various BRDF parameterization methods. Based on this, we select a specular-glossiness (Spec-Gloss) [16] material parameterization. We find that this decomposition scheme extracts reasonable material attributes more effectively compared to other rendering models. Instead of 3DGS, we adapt this rendering model for a surfel-based scene representation [8], enabling high-quality reconstruction by providing high accuracy surface modeling. We leverage priors from diffusion models to steer the optimization of geometry and materials and to reduce ambiguity in the inverse solving process, resulting in more realistic material attributes and environment lighting. We propose an efficient environment map training strategy that, through a coarse-to-fine approach, accelerates scene reconstruction, enables high-resolution environment map training, and enhances reconstruction quality while assisting the optimization of object material properties. Experimental results demonstrate that our method achieves high-quality geometry and material reconstruction, while delivering more photorealistic relighting results, significantly surpassing those of previous approaches. In summary, our contributions include:
@@ 2 Related Work
Figure 2: Overview of our framework built on 2DGS [8]. Gaussian splats rasterize to a G buffer of albedo, roughness, $F_{0}$, indirect color, and surface normals. A differentiable prefiltered environment cubemap with mipmaps provides lighting in a physically based deferred renderer. The HDR environment map is learned in a coarse-to-fine manner. Supervision uses an sRGB photometric loss between shaded output and ground truth (GT), plus normal and diffuse priors that reduce ambiguity between geometry, materials, and lighting.
@@ 2.1 Neural Rendering for Glossy Scenes
Reconstructing glossy scenes is hard because specular cues are high frequency and view dependent, which creates multi-view inconsistencies and entangles geometry, materials, and light. NeRF [23], DVGO [26], and NeuS [32] advance geometry and novel view synthesis but struggle with specular fidelity or slow training. Ref-NeRF [30] adds a simplified BRDF and implicitly learns reflected radiance, but remains slow and prevents explicit relighting. 3DGS [11] improves efficiency and quality but is not relightable and has no explicit surface normals in the representation. GaussianShader [9] extends 3DGS with a simple specular term, yet yields noisy normals, which hurts geometry, materials, and lighting recovery. R3DG [6] improves geometry by attaching BRDF to points and using ray-based visibility, but per-point queries scale poorly on dense point clouds. 3DGS-DR [12] scales better and adopts deferred shading with normal propagation from the shaded buffer to the primitives, which stabilizes geometry, but it does not use a full microfacet BRDF, so disentanglement and relighting fidelity remain limited. 2DGS [8] introduces view-consistent planar disks with explicit normals, better suited to reconstructing glossy objects. Subsequent works [37, 33, 15, 27] adopt 2DGS with pixel-level deferred shading and differentiable environment lighting, alleviating geometry issues and scaling constraints. We build on this line of work and use 2DGS with deferred shading as our foundation.
@@ 2.2 Inverse Rendering
Inverse rendering seeks materials and illumination that explain multi-view images. Early neural methods incorporated BRDFs and environment lighting into NeRF-style pipelines but were slow, unable to relight, or required prior geometry or lighting [30, 13, 36, 24]. SDF-based approaches improved geometry and material quality but were costly [19, 21, 5, 7]. Gaussian splats with forward shading blurred specular cues and hindered disentanglement [9, 6], while deferred pipelines separated rasterization from shading and improved stability [12]. Recent 2DGS work coupled deferred shading with image-based lighting, preserving sharp highlights and improving optimization [37, 33, 27, 15].
Material parameterization and BRDF remain central. Disney’s metallic-roughness is intuitive but ties specular reflectance to base color, harming identifiability [3]. Specular-glossiness treats specular reflectance as free, alleviating entanglement. Hybrid parameterizations [9, 33] introduce specular tint but are incompatible with standard graphics workflows. ReCap [18] adopts Spec-Gloss in 3DGS but needs multiple illuminations. We also adopt Spec-Gloss in a split-sum approximation with 2DGS and deferred shading for better surface reconstruction, incorporating diffuse color and normal priors. Unlike [12, 33], we avoid bounding positive HDR radiance, using negative-only clipping to preserve bright highlights and improve material-lighting disentanglement. To reduce training time, we propose coarse-to-fine cubemap upsampling paralleling frequency-progressive spherical-harmonics lighting [20].
@@ 2.3 Inverse Rendering Priors
As already established, inverse rendering is ill-posed since many material–lighting pairs can explain the same images, which harms downstream scene editing tasks. To reduce this ambiguity, prior work adds structure in different ways, such as constraining illumination [24], using data-driven BRDF priors [36, 19], supplying pretrained or externally supervised geometry [21, 5], seeding materials with predictors [4], or injecting surface-normal and depth priors during optimization [27]. We follow this line by using learned normal and diffuse color priors through the diffusion models StableNormal [34] and StableDelight [25] for early stabilization and to steer the optimization in the right direction.
@@ 3 Methodology
Our method integrates the Specular-Glossiness BRDF rendering model with a surfel-based scene representation and a pixel-level deferred shading technique (§ 3.2). To balance speed and quality, we progressively upsample the environment light cubemap in a coarse-to-fine manner (§ 3.3). During scene reconstruction, we leverage diffusion priors on normals and material properties to reduce ambiguities and better disentangle materials from illumination (§ 3.4). The proposed method is shown in Fig. 2.
@@ 3.1 Preliminary
Surfel-based Scene Representation.The main limitations of 3DGS [11] are its multi-view geometric inconsistencies and the absence of explicit surface normals, both of which hinder inverse rendering of glossy objects. To address this, we build on 2DGS [8], which collapses 3D Gaussians into view-consistent planar disks and exposes explicit surface normals aligned with their shortest axis. This design yields thinner surfaces and enables more stable optimization, making it better suited for high-frequency specular effects.
A 2D splat is defined by its center $\mathbf{p}_{k}$, two principal tangential vectors $\mathbf{t}_{u},\mathbf{t}_{v}$, a scaling vector $S=(s_{u},s_{v})$. The orientation of the 2D primitive is given by a $3\times 3$ rotation matrix $\mathcal{R}=[\mathbf{t}_{u},\mathbf{t}_{v},\mathbf{t}_{w}]$ and its explicit normal is defined as $\mathbf{n}=\mathbf{t}_{u}\times\mathbf{t}_{v}$. The 2D Gaussian value of a point $\mathbf{u}=(u,v)$ in the $uv$ local tangent space can be determined via $\mathcal{G}(u,v)=exp(-(u^{2}+v^{2})/2)$. The position $p_{k}$, scaling $S$, and rotation $R$ are learnable sets of parameters and combined with a learnable opacity value $\alpha$, can be used to rasterize various properties such as color by blending ordered points overlapping the pixel:
@@
where $c$ refers to any Gaussian-level parameter, and $\mathcal{\hat{G}}$ is a low-pass-filtered $\mathcal{G}$. More details can be found in [8].
@@ 3.2 BRDF Parameterization and Rendering
The widely used Disney BRDF [3] employs a metallic–roughness (MR) parameterization with base color $b\in\mathbb{R}^{3}$, metallic $m\in[0,1]$, and roughness $r\in[0,1]$. The specular reflectance is implicitly defined via:
@@
In contrast, the specular-glossiness (SG) parameterization decouples diffuse and specular components by explicitly learning the specular reflectance $F_{0}$ as a free RGB parameter. This enables a clean separation of appearance into diffuse albedo $k_{d}$, specular reflectance $F_{0}$, and roughness $r$, improving the identifiability of material properties, especially under glossy reflections.
@@
IBL with Split-Sum Approximation. To render the interaction of light with glossy surfaces, we adopt the split-sum approximation of image-based lighting (IBL) [10]. This physically based formulation separates illumination into diffuse and specular terms and enables efficient approximation of the rendering equation using precomputed lookup tables and mipmapped environment maps.
$L_{d}(\mathbf{n})=\int_{\Omega}\tfrac{1}{\pi}\max(\mathbf{n}\cdot\mathbf{\omega}_{i},0)E(\mathbf{\omega}_{i})d\mathbf{\omega}_{i}.$ (3)
where $\mathbf{n}$ is the surface normal obtained from surfel-based scene representation, $\mathbf{\omega}_{i}$ is a light direction, and $E$ is the environment radiance. The specular term is approximated by prefiltering the environment map with GGX importance sampling [28] and combining the result with a 2D BRDF lookup table ($BRDF_{LUT}$):
$L_{s}(\mathbf{v},\mathbf{n},r)\approx\text{PrefilterEnv}(\mathbf{\omega}_{r},r)\odot\text{BRDF}_{\text{LUT}}(r,\mathbf{n}\cdot\mathbf{v}),$ (4)
where $\mathbf{v}$ is the viewing direction, $\mathbf{\omega}_{r}$ is the reflection direction around the surface normal, and $r$ is the roughness. $\text{PrefilterEnv}(\cdot)$ is a pre-filtering function for the split-sum approximation. The split-sum approximation enables efficient and stable inverse rendering by precomputing the BRDF integration into a LUT and using a prefiltered environment map, avoiding costly high-dimensional Monte Carlo sampling. The BRDF lookup-table $\text{BRDF}_{\text{LUT}}$ provides the Fresnel coefficients $\beta_{1}$, $\beta_{2}$ used in the computation of outgoing color $L_{o}$:
$L_{o}=k_{d}L_{d}+(F_{0}\beta_{1}+\beta_{2})L_{s}.$ (5)
Figure 3: Relighting of bell and tbell scenes from the Glossy Synthetic dataset [21] with different material parameterizations integrated into our methodology, demonstrating the improved material-lighting disentanglement of the Spec-Gloss variant.
Deferred Shading. Similar to [12, 27, 33], we adopt pixel-level deferred shading and alpha-blend the per-Gaussian attributes—albedo $\mathbf{k}_{d,i}$, roughness $r_{i}$, specular reflectance $\mathbf{F}_{0,i}$, and surface normals $\mathbf{n}_{i}$—using Eq. 1 to construct the screens-pace G-buffer with dense per-pixel maps $K_{d}$, $R$, $F_{0}$, and $N$. Physically-based BRDF shading is then applied at the pixel level. This decoupling of geometry and material estimation from shading leads to sharper highlights, better supervision signals during training, and robustness to view-dependent specularities compared to forward rendering.

### v2

# Spec-Gloss Surfels and Normal-Diffuse Priors for Relightable Glossy Objects

@@ Abstract
Accurate reconstruction and relighting of glossy objects remains a longstanding challenge, as object shape, material properties, and illumination are inherently difficult to disentangle. Existing neural rendering approaches often rely on simplified BRDF models or parameterizations that couple diffuse and specular components, which restrict faithful material recovery and limit relighting fidelity. We propose a relightable framework that integrates a microfacet BRDF with the specular-glossiness parameterization into 2D Gaussian Splatting with deferred shading. This formulation enables more physically consistent material decomposition, while diffusion-based priors for surface normals and diffuse color guide early-stage optimization and mitigate ambiguity. A coarse-to-fine environment map optimization accelerates convergence, and negative-only environment map clipping preserves high-dynamic-range specular reflections. Extensive experiments on complex, glossy scenes demonstrate that our method achieves high-quality geometry and material reconstruction, delivering substantially more realistic and consistent relighting under novel illumination compared to existing Gaussian splatting methods. The source code is available at https://github.com/gkouros/SpecGloss-GS.
@@ 1 Introduction
Incorporating real-world objects into virtual environments is a fundamental task in applications such as digital twins and virtual reality. A major challenge arises in accurately reconstructing objects from the physical world and reproducing realistic lighting effects within specific virtual settings. This difficulty is particularly pronounced for highly reflective objects, where the lack of multi-view consistency renders traditional approaches, such as multi-view stereo (MVS), ineffective and makes reliable reconstruction a non-trivial problem.
@@
The choice of scene representation and rendering model introduces unique trade-offs in performance. For instance, approaches based on Radiance Fields [mildenhall2020nerf, SunSC22dvgo, barron2022mip360] or 3D Gaussian Splatting (3DGS) [kerbl20233dgs] often fail to accurately capture object surfaces, thereby limiting their ability to model surface reflections. In contrast, methods employing Signed Distance Functions (SDFs) [wang2021neus, ge2023ref, liu2023nero, liang2023envidr] provide precise surface geometry, but their high computational cost poses challenges for real-time rendering. Another critical component of inverse rendering is the rendering model, which decomposes object materials into multiple attributes and employs predefined functions to synthesize appearance from target viewpoints. To achieve differentiability, various methods approximate the rendering equation integrals differently. For example, Ref-NeRF [verbin2022refnerf] uses an Integrated Directional Encoding, GaussianShader [jiang2024gshader] employs specular GGX [Walter2007Microfacet], Ref-GS [zhang2024refgs] adopts a spherical-mip encoding, and 3DGS-DR [ye2024gsdr] leverages deferred rendering. While these diverse formulations yield varying reconstruction quality, they share a fundamental limitation: the inverse solving process suffers from an inherent ambiguity, where multiple material-lighting combinations can produce identical appearances. Consequently, reconstructed material attributes and recovered lighting often deviate from ground truth values, leading to degraded rendering quality under novel lighting conditions. This is not a new observation, but a long-standing open problem going back to early work on intrinsic images [BarrowTenenbaum1978], highlighted more recently in [Kouros_2024_CVPR].
@@
In this work, we reduce ambiguity and improve relighting quality. We start by investigating the relationship between rendering models and relighting effects. Specifically, we analyze the coupling between object materials and various BRDF parameterization methods. Based on this, we select a specular-glossiness (Spec-Gloss) [Lagarde2014MovingPBR] material parameterization. We find that this decomposition scheme extracts reasonable material attributes more effectively compared to other rendering models. Instead of 3DGS, we adapt this rendering model for a 2DGS-based scene representation [huang20242dgs], enabling high-quality reconstruction by providing high accuracy surface modeling. We leverage priors from diffusion models to steer the optimization of geometry and materials and to reduce ambiguity in the inverse solving process, resulting in more realistic material attributes and environment lighting. We propose an efficient environment map training strategy that, through a coarse-to-fine approach, accelerates scene reconstruction, enables high-resolution environment map training, and enhances reconstruction quality while assisting the optimization of object material properties. Experimental results demonstrate that our method achieves high-quality geometry and material reconstruction, while delivering more photorealistic relighting results, significantly surpassing those of previous approaches. In summary, our contributions include:
@@ 2 Related Work
Figure 2: Overview of our framework built on 2DGS [huang20242dgs]. Gaussian splats rasterize to a G buffer of albedo, roughness, $F_{0}$, indirect color, and surface normals. A differentiable prefiltered environment cubemap with mipmaps provides lighting in a physically based deferred renderer. The HDR environment map is learned in a coarse-to-fine manner. Supervision uses an sRGB photometric loss between shaded output and ground truth (GT), plus normal and diffuse priors that reduce ambiguity between geometry, materials, and lighting.
@@ 2.1 Neural Rendering for Glossy Scenes
Reconstructing glossy scenes is hard because specular cues are high frequency and view dependent, which creates multi-view inconsistencies and entangles geometry, materials, and light. NeRF [mildenhall2021nerf], DVGO [SunSC22dvgo], and NeuS [wang2021neus] advance geometry and novel view synthesis but struggle with specular fidelity or slow training. Ref-NeRF [verbin2022refnerf] adds a simplified BRDF and implicitly learns reflected radiance, but remains slow and prevents explicit relighting. 3DGS [kerbl20233dgs] improves efficiency and quality but is not relightable and has no explicit surface normals in the representation. GaussianShader [jiang2024gshader] extends 3DGS with a simple specular term, yet yields noisy normals, which hurts geometry, materials, and lighting recovery. R3DG [R3DG2024] improves geometry by attaching BRDF to points and using ray-based visibility, but per-point queries scale poorly on dense point clouds. 3DGS-DR [ye2024gsdr] scales better and adopts deferred shading with normal propagation from the shaded buffer to the primitives, which stabilizes geometry, but it does not use a full microfacet BRDF, so disentanglement and relighting fidelity remain limited. 2DGS [huang20242dgs] introduces view-consistent planar disks with explicit normals, better suited to reconstructing glossy objects. Subsequent works [zhang2024refgs, yao2025refgaussian, kouros2025rgsdr, tong2025gs2dgs] adopt 2DGS with pixel-level deferred shading and differentiable environment lighting, alleviating geometry issues and scaling constraints. We build on this line of work and use 2DGS with deferred shading as our foundation.
@@ 2.2 Inverse Rendering
Inverse rendering seeks materials and illumination that explain multi-view images. Early neural methods incorporated BRDFs and environment lighting into NeRF-style pipelines but were slow, unable to relight, or required prior geometry or lighting [verbin2022refnerf, kouros2023refdvgo, zhang2021nerfactor, srinivsan2021nerv]. SDF-based approaches improved geometry and material quality but were costly [liang2023envidr, liu2023nero, fan2023factoredneus, ge2023ref]. Gaussian splats with forward shading blurred specular cues and hindered disentanglement [jiang2024gshader, R3DG2024], while deferred pipelines separated rasterization from shading and improved stability [ye2024gsdr]. Recent 2DGS work coupled deferred shading with image-based lighting, preserving sharp highlights and improving optimization [zhang2024refgs, yao2025refgaussian, tong2025gs2dgs, kouros2025rgsdr].
Material parameterization and BRDF remain central. Disney’s metallic-roughness is intuitive but ties specular reflectance to base color, harming identifiability [burley2012disney]. Specular-glossiness treats specular reflectance as free, alleviating entanglement. Hybrid parameterizations [jiang2024gshader, yao2025refgaussian] introduce specular tint but are incompatible with standard graphics workflows. ReCap [li2025recap] adopts Spec-Gloss in 3DGS but needs multiple illuminations. We also adopt Spec-Gloss in a split-sum approximation with 2DGS and deferred shading for better surface reconstruction, incorporating diffuse color and normal priors. Unlike [ye2024gsdr, yao2025refgaussian], we avoid bounding positive HDR radiance, using negative-only clipping to preserve bright highlights and improve material-lighting disentanglement. To reduce training time, we propose coarse-to-fine cubemap upsampling paralleling frequency-progressive spherical-harmonics lighting [liu2017progressivesh].
@@ 2.3 Inverse Rendering Priors
As already established, inverse rendering is ill-posed since many material–lighting pairs can explain the same images, which harms downstream scene editing tasks. To reduce this ambiguity, prior work adds structure in different ways, such as constraining illumination [srinivsan2021nerv], using data-driven BRDF priors [zhang2021nerfactor, liang2023envidr], supplying pretrained or externally supervised geometry [liu2023nero, fan2023factoredneus], seeding materials with predictors [du2024gsid], or injecting surface-normal and depth priors during optimization [tong2025gs2dgs]. We follow this line by using learned normal and diffuse color priors through the diffusion models StableNormal [ye2024stablenormal] and StableDelight [stabledelight2025] for early stabilization and to steer the optimization in the right direction.
@@ 3 Methodology
Our method integrates the Specular-Glossiness BRDF rendering model with a 2DGS-based scene representation and a pixel-level deferred shading technique (§ 3.2). To balance speed and quality, we progressively upsample the environment light cubemap in a coarse-to-fine manner (§ 3.3). During scene reconstruction, we leverage diffusion priors on normals and material properties to reduce ambiguities and better disentangle materials from illumination (§ 3.4). The proposed methodology is illustrated in Fig. 2.
@@ 3.1 Preliminary
Surfel-based Scene Representation. The main limitations of 3DGS [kerbl20233dgs] are its multi-view geometric inconsistencies and the absence of explicit surface normals, both of which hinder inverse rendering of glossy objects. To address this, we build on 2DGS [huang20242dgs], which collapses 3D Gaussians into view-consistent planar disks and exposes explicit surface normals aligned with their shortest axis. This design yields thinner surfaces and enables more stable optimization, making it better suited for high-frequency specular effects.
A 2D splat is defined by its center $\mathbf{p}_{k}$, two principal tangential vectors $\mathbf{t}_{u},\mathbf{t}_{v}$, a scaling vector $S=(s_{u},s_{v})$. The orientation of the 2D primitive is given by a $3\times 3$ rotation matrix $\mathcal{R}=[\mathbf{t}_{u},\mathbf{t}_{v},\mathbf{t}_{w}]$ and its explicit normal is defined as $\mathbf{n}=\mathbf{t}_{u}\times\mathbf{t}_{v}$. The 2D Gaussian value of a point $\mathbf{u}=(u,v)$ in the uvuv local tangent space can be determined via $\mathcal{G}(u,v)=exp(-(u^{2}+v^{2})/2)$. The position $p_{k}$, scaling $S$, and rotation $R$ are learnable sets of parameters and combined with a learnable opacity value $\alpha$, can be used to rasterize various properties such as color by blending ordered points overlapping the pixel:
@@
where $c$ refers to any Gaussian-level parameter, and $\mathcal{\hat{G}}$ is a low-pass-filtered $\mathcal{G}$. More details can be found in [huang20242dgs].
@@ 3.2 BRDF Parameterization and Rendering
The widely used Disney BRDF [burley2012disney] employs a metallic–roughness (MR) parameterization with base color $b\in\mathbb{R}^{3}$, metallic $m\in[0,1]$, and roughness $r\in[0,1]$. The specular reflectance is implicitly defined via:
@@
In contrast, the specular-glossiness (SG) parameterization decouples diffuse and specular components by explicitly learning the specular reflectance $F_{0}$ as a free RGB parameter. This enables a clean separation of appearance into diffuse albedo $\rho_{d}$, specular reflectance at normal incidence $F_{0}$, and roughness $r$, improving the identifiability of material properties, especially under glossy reflections.
@@
IBL with Split-Sum Approximation. To render glossy surfaces, we adopt the split-sum approximation for image-based lighting (IBL) [Karis2013ibl], which decomposes illumination into diffuse and specular terms. The diffuse term $L_{d}(\mathbf{n})$ is obtained by sampling a low-frequency irradiance map, $L_{d}(\mathbf{n})\approx\text{PrefilterEnv}(\mathbf{n},r^{\max})$, where $\text{PrefilterEnv}(\cdot,r)$ samples a mipmapped environment map prefiltered over roughness $r$ and $r^{\max}$ denotes the maximum roughness (coarsest mip level). For the specular term, we prefilter the environment map with GGX importance sampling [trowbridge1975average], yielding a direct specular IBL component $L_{\text{dir}}(\mathbf{v},\mathbf{n},r)\approx\text{PrefilterEnv}(\boldsymbol{\omega}_{r},r)$, where $\mathbf{v}$ is the viewing direction, $\boldsymbol{\omega}_{r}$ is the reflection direction about the surface normal $\mathbf{n}$, and $r$ is the roughness. Indirect specular interreflections are modeled by a low-frequency spherical-harmonics term $L_{\text{ind}}(\boldsymbol{\omega}_{r})$ (see supplementary). A visibility term $V\in\{0,1\}$ blends direct and indirect specular illumination:
$L_{s}(\mathbf{v},\mathbf{n},r)=V\,L_{\text{dir}}(\mathbf{v},\mathbf{n},r)+(1-V)\,L_{\text{ind}}(\boldsymbol{\omega}_{r}).$ (3)
Following the split-sum formulation, the remaining GGX BRDF integration is precomputed in a 2D lookup table $\text{BRDF}_{\text{LUT}}(r,\mathbf{n}\cdot\mathbf{v})$, which returns scale and bias coefficients $\beta_{1}$ and $\beta_{2}$ for the Fresnel term. The outgoing radiance is then
$L_{o}(\mathbf{v},\mathbf{n})=\rho_{d}\,L_{d}(\mathbf{n})+\bigl(F_{0}\,\beta_{1}+\beta_{2}\bigr)\odot L_{s}(\mathbf{v},\mathbf{n},r),$ (4)
where $\rho_{d}$ is the diffuse albedo, $F_{0}$ is the specular reflectance at normal incidence, and $\odot$ denotes element-wise multiplication.
Figure 3: Relighting of bell and tbell scenes from the Glossy Synthetic dataset [liu2023nero] with different material parameterizations integrated into our methodology, demonstrating the improved material-lighting disentanglement of the Spec-Gloss variant.
Deferred Shading. Similar to [ye2024gsdr, tong2025gs2dgs, yao2025refgaussian], we adopt pixel-level deferred shading and alpha-blend the per-Gaussian attributes (albedo $\rho_{d,i}$, roughness $r_{i}$, specular reflectance $F_{0,i}$, and surface normals $\mathbf{n}_{i}$) using Eq. 1 to construct the screen-space G-buffer with dense per-pixel maps $P_{d}$, $R$, $F_{0}$, and NN. Physically-based BRDF shading is then applied at the pixel level. This decoupling of geometry and material estimation from shading leads to sharper highlights, better supervision signals during training, and robustness to view-dependent specularities compared to forward rendering.